answer question 3and 4Question 2 (2 points) If 0.25 g of Na2HPO4 (disodium phosphate) is added to the 50 mL 0.30 M KH2PO4 solution, calculate the p
Answer #1

3)

pKb of NH3 = 4.74

moles of NH3 = 50.0 x 1.0 / 1000 = 0.050 mol

pH = 9.00

pOH = 6.00

pOH = pKb + log [NH4Cl / NH3]

6.00 = 4.74 + log [NH4Cl / 0.050]

moles of NH4Cl = 0.9098

0.50 x V = 0.9098

V = 1.82

volume of NH4Cl = 1.82 L

4)

mmoles of NaOH = 10 x 0.24 = 2.4

mmoles of CH3COOH = 50 x 0.20 = 10

CH3COOH +   NaOH   - ----------> CH3COO- +   H2O

    10                  2.4                              0                   0

    7.6                   0                             2.4

pH = pKa + log [CH3COO- / CH3COOH]

     = 4.74 + log [2.4 / 7.6]

pH = 4.24